Sophie wants to see if there is evidence that average lead level for a lake is lowerthan 15 ppm. After she collected water samples from randomly selected 25 spots of the lake, the samples are tested and data are collected. When she analyzed the data, she calculated her test statistic t = - 2.25. Which of the following is correct based on the information, if the significance level is set as 0.02?

Group of answer choices

P-value < 0.02, and there is evidence that the lead level for the lake is lower than 15 ppm.

P-value > 0.02, and there is not enough evidence that the lead level for the lake is lower than 15 ppm.

P-value > 0.02, and there is evidence that the lead level for the lake is lower than 15 ppm.

P-value < 0.02, and there is not enough evidence that the lead level for the lake is lower than 15 ppm.
Answer #1

This is the left tailed test .

df = 25 - 1 = 24

P-value = P(t 24 < -2.25) = 0.0169

P-value < 0.02, and there is evidence that the lead level for the lake is lower than 15 ppm.